OpenStudy (anonymous):

express in terms of i √-50=

OpenStudy (anonymous):

(50)^1/2i

OpenStudy (anonymous):

\[\sqrt{50}*i\]

OpenStudy (anonymous):

a little further: \[\sqrt{-50} = \sqrt{-1}*\sqrt{50} = i\sqrt{25*2} = i\sqrt{25}*\sqrt{2} = 5i\sqrt{2}\]